# Embodying Control in Soft Multistable Grippers from morphofunctional co-design

### v1

@@ Abstract
Soft robots are distinguished by their ability to adapt to various environments and interaction forces through shape-changing and responsiveness to external stimuli. Their flexible materials provide inherent safety, allowing them to perform tasks that are nearly impossible for rigid robots. However, controlling the configuration of soft robots is challenging due to their highly nonlinear material response and infinite deflection degrees of freedom, often necessitating sophisticated sensors and complex closed-loop control systems. One potential solution is to discretize the infinite-dimensional configuration space of soft robots into a finite but sufficiently large number of functional shapes. This approach facilitates modeling and control without sacrificing compliance and adaptability. This study explores a co-design strategy for a class of pneumatically actuated soft grippers with multiple encoded stable states, enabling desired functional shape and stiffness reconfiguration. Combining a PneuNet bending actuator with dome-shaped bistable units allowed us to encode stable configurations and design them to exhibit desired kinematic and stiffness responses. This architecture enables position and force regulation of soft robots without closed-loop control. We present an energy-based analytical model for these soft multistable grippers based on a lattice of lumped parameter elements. This model maps the soft robots’ infinite-dimensional configuration space into discrete stable states, allowing us to predict the system’s final state and the dynamic behavior between each state upon actuation. Our approach introduces a general method to capture the soft robots’ response with the lattice lumped parameters using automatic relevance determination regression, facilitating inverse co-design. The resulting computationally efficient model enables us to explore the configuration space in a tractable manner, allowing the inverse co-design of our robots by setting desired targeted positions with optimized stiffness of the set targets. Our model is validated against FE simulations and experimental tests for different target points, demonstrating the potential of discretizing the modeling and design of soft robots using multistable structures. We showcase the utility of our modeling framework through inverse design for various applications and tasks, combining diverse mechanical behaviors into the same soft robot. Our strategy of mapping the modeling and co-design of soft grippers into a discretized space offers a framework for controlling soft robots by exploiting the nonlinear mechanics of multistable structures, thus opening avenues for embodying mechanical intelligence into soft structures.
@@ 1 Introduction
Soft robots have gained interest due to their ability to interact with their environment, adapt to external stimuli, and protect against external disturbances [1, 2, 3]. The inherent safety provided by the characteristic low modulus of materials used in soft robots allows them to perform nearly impossible tasks for their rigid counterpart[4]. The interplay between soft mechanics and controls intrinsic to soft robotics gives rise to innovative solutions for tasks ranging from simple grasping to complex manipulation and interaction[5, 6, 7].However, this interplay also poses challenges complicating their modeling and control. These challenges include their infinite dimensionality, material nonlinearity, and large deformations that most soft robots exhibit[8]. As a result, sensory systems[9] and complex models/algorithms [10, 11] are required to implement closed-loop control, leading to computationally demanding models to represent the robot’s behavior. Embedding sensing and control in the robots’ architecture [12] offers a new strategy to address some of the main problems of soft robotics reducing the use of complex algorithms and models to predict the robot’s behavior. For example, granular material-based universal grippers [13] can adapt to different object shapes and grasp them without the need for closed-loop control based on their mechanical response. Zoe et al. [14] developed a methodology to use various soft actuators as pneumatic sensors, enabling distance sensing, shape sensing, profile scanning, and stiffness determination. By incorporating the robot’s geometry and body into the design, these examples illustrate the possibility of closing the control loop to realize desired tasks by the intelligent co-design of the robot’s morphology and functional behavior.
@@
Multistable structures offer an alternative path to achieve control of soft robotics with feedback via the programming of input-specific stable and defined shapes[15, 16, 17]. These types of structures display more than one energetically favorable configuration, giving the possibility to reach different final shapes from specific actuation inputs. They are often the result of assembling multiple classical bistable substructures including constrained beams/trusses [18, 19], constrained dielectric elastomers [20], shells [21, 22, 23, 24], compliant mechanisms [25], or inflatable structures [26]. Combining multistable structures with soft robots can reduce the number of degrees of freedom and complexity to achieve control in soft robotic devices, ultimately enhancing their functionality by providing the ability to reach desired set points without closed-loop feedback[27, 28, 29]. Presently, a wide variety of soft machines leverage mechanical instabilities to improve the design and performance of their soft machines [30, 31, 32, 33] by programming different stable configurations and utilizing the rapid energy release during snap-through instability [34, 35, 36, 37]. Conrad et al. show an electronics-free controller for a complex pneumatic system [38] based on soft logic gates that can be incorporated into different soft machines. Similarly, Peretz et al. utilized mechanical instabilities to create fluidic logic and reproduce valving functionality using entirely soft elements[39]. Melancon et al. [40] used a bistable origami pattern to program different target points to its actuator, achieved by a single pressure source and predefined pressure path. Van Raemdonck et al. [41] geometrically tuned an actuator to generate complex actuation sequences from a single input by delaying the unit snap-thought. Moreover, by leveraging the stable points in the structure, the robot’s final state can be predicted based solely on its structural response, thereby defining set points in the mechanical behavior of the structure. This approach can simplify the complexity of predictive or data-driven algorithms currently needed to design and control soft robots, ultimately making soft robots more accessible and practical. However, the full potential of using multistability for soft robotics can only be realized by establishing computationally efficient methodologies for their morphological and functional co-design that circumvent the long runtime roadblocks imposed by the predominant use of conventional Finite Element (FE) packages to predict the geometry and response characteristics associated with the desired stable states.
@@
We introduce a framework for the modeling, control, and inverse design of soft robotics by mapping into a discretized space the desired kinematic configurations leveraging multistability (see Fig. 1). We achieve this by incorporating different dome-shape programmable units addressable via pneumatic actuation featuring bistable, pseudo-bistable, and monostable [21, 42] mechanical behaviors (see Fig. S2). As a result, different set points in the form of specific kinematical shapes are embodied into the system’s response (see Fig 1), giving our soft robots the inherent capability of reaching different positions with zero error. This characteristic allows for robust soft robot design, as the bistable units provide intrinsic kinematic control dictated by the embodied mechanical response. The position control inherent to the designed multistable structure can serve as set points for triggering different control actions. The discretized response afforded by the proposed topology enables the introduction of an energy-based analytical modeling framework, allowing the efficient shape and structural property prediction of a multistable soft robot appendage topology (Dome Phalanx Finger - DPF) that ultimately enables the inverse co-design of the robot’s functionality and control via the morphology. Our modeling approach is based on lumped parameter elements that can capture different mechanical behaviors of the dome unit via mechanics-informed identification, mapping the geometry of our robot to its mechanical response. The collective interaction of these elastic elements dictates the soft robot’s configuration, stiffness, and dynamic response. We demonstrate the versatility of our approach by performing the inverse co-design of different DPF topologies, giving specific target positions and geometrical constraints while maximizing the programmed states’ stiffnesses. Finally, we show that the time response can be incorporated into the DPF morphology by leveraging the viscous effects of the material, giving a new avenue for programming this class of robots’ time response, illustrated through pick-and-place applications. The presented modeling and inverse co-design approach leverages geometry and the resulting structural response to encode an electronics-free form of mechanical intelligence that simplifies the actuation and control of soft robots with implications on their accessibility, recyclability, and cost.
@@
Figure 1: Multistable Soft Robot with Pre-programmed Position and Stiffness. The four-finger robot demonstrates four grasping set points and embodied control. The geometry and mechanical behavior of the robot encode the input and output behaviors in the energy landscape (G(s)). Feedback H(s) is derived from the dynamic response of the robot, which yields different attractors for each stable set point in the robot. The system’s output shows four set points as static minima in the energy landscape and attractors in the phase portrait. Set points (energy minima) are accessed by activating the dome units: State 1 $\rightarrow$ Stress-free state and State 2 $\rightarrow$ Inverted state.
@@ 2.1 Embodied position and stiffness control
We start by incorporating dome-shaped units into the PneuNet bending actuator topology[43] (see Fig. S1) to create our Dome Phalanx Finger (DPF) topology (Fig. 2a). The finger derives its behavior from the mechanical response of the dome-shaped shell elements and the interaction between each segment. As the units invert, the top section of the finger expands, while the limiting layer (see Fig. S1d) retains its original length. This results in global curvature and different final stable states, depending on the number of active units (see Fig. S2c).
@@
These dome-shaped units program different mechanical responses and stable states into the system, effectively discretizing the infinite-dimensional deformation space into a manageable number of kinematic configurations dependent on the unit cell geometry, each attainable via well-defined open-loop inputs. The mechanical response and stable configurations of the DPF can be tuned by adjusting the dome height per thickness ($H_{i}$), dome thickness ($t$), number of segments, unit cell size (UC), limiting layer thickness ($t_{lim}$), unit cell length ($U_{L}$), air chamber dimensions ($W_{c}$ and $L_{c}$), and the spacing between adjacent cells ($U_{sep}$) (see Fig. 2a and Fig. S1a for reference). The dome unit geometric parameters ($H$ and $t$) determine whether the mechanical response is monostable (Fig. S1c (i)), metastable (Fig. S1c (ii)), or bistable (Fig. S1c (iii)). The other geometrical parameters influence the interaction between neighboring segments without altering the unit’s stability, while, affecting the global curvature.
@@
Given the broad range of geometrical configurations, each finger can be geometrically tuned to exhibit different mechanical responses: a bistable response, where the final shape is retained after dome inversion[21, 22, 44] (Fig. 2b $H_{i}$ = 4 mm and $H_{i}$ = 5 mm); a pseudo-bistable (metastable) response, where each unit undergoes a snap-through instability and the shape is retained for a time $t=\tau$, before returning to its stress-free state from the combined effects of the unit’s viscoelastic material behavior and dome geometry [34, 45, 46] (Fig. 2b, H = 3 mm); and finally, a monostable response, where the material behavior dominates over the geometric response and the finger responds similarly to a PneuNet bending actuator.
@@
Figure 2: Dome Phalanx Finger (DPF) with different encoded mechanical responses and behaviors. a) 5 segment DPF geometry parameters. ($i=5$) b) Effect of dome height on finger response (Every segment with the same height). $H_{i}$ = 5 mm and $H_{i}$ = 4 mm show a bistable behavior, and $H_{i}$ = 3 mm shows a pseudo-bistable response (self-resetting system). c) Time lapse of two gripper architectures showing both bistable and pseudo-bistable responses.
@@
These unique properties can be further exploited in a gripper architecture. By changing the dome height, one can program a gripper composed of two fingers to pick and hold an object after actuation (Fig. 2c, H = 4 mm) or pick an object and release it as the units reset (Fig. 2c, H = 3 mm), the latter of which is ideal for pick-and-place applications typically performed by conventional soft grippers. The number of stable states that can be encoded using the DPF morphology equals that of the number of constitutive units where each unit is tunable independently. Consequently, different unit geometries can be combined to achieve multiple stable states or combined behaviors within the same overall geometry. This allows for different energy minima and time responses to be embodied within the same finger, which can be accessed by a single pressure input and serve as discrete and functional robot configurations. To effectively incorporate, combine, and predict the finger’s behavior based on its geometry, a simple yet reliable modeling framework is necessary. Given the ample design space and the computational cost of using finite element (FE) models for our finger topology, we propose an energy-based spring model that captures the complex mechanical behaviors needed to design multistable soft robots with desired functional characteristics.
@@ 2.2 Model derivation and mechanics-informed parameter determination
Establishing an analytical model to capture the global behavior of the DPF requires consideration of both the unit cells’ and the global geometric parameters. By capturing the contribution of each constitutive unit of the DPF and their interactions, we can construct an energy landscape with a simplified model, where each of the minima corresponds to a stable state and a final position of the system. In this way, the state of the system ($q$) can be extracted from the resulting strain energy potential via a minimization process, which dictates the programmed stable shapes with their geometrical and stiffness characteristics.
Figure 3: An energy-based model for static and dynamic analysis of multistable soft robots based on the DPF topology. a) Spring lattice model with linear, rotational, and nonlinear springs. b) Static stable states predicted by the lattice model for $H_{1}$ = $H_{2}$ = 4 mm, $H_{3}$ = $H_{4}$ = 4.5 mm, and $H_{5}$ = 5 mm. c) Dynamic response for the soft robot with three bistable units and two metastable units ($H_{1}=4$ mm, $H_{2}=4.5$ mm, $H_{3}=5$ mm and $H_{4}=H_{5}=3$ mm). The model captures the snap-through of each bistable unit and the snap-back of the pseudo-bistable units.
The DPF geometry is represented by a lattice array comprising nonlinear, linear, and torsional springs for the static response and point masses and dash-pots for the dynamic response (Fig. 3a). The energy for the springs used in the lattice can be written as follows[47].
Nonlinear: $\text{{Nonlinear:} }E_{NL}=\frac{1}{2}k_{b}x^{2}\left(1+\left((1-\alpha)\left(\left(\frac{x}{d}\right)^{2}-2\frac{x}{d}\right)\right)\right)$ (1)
Linear: $\text{{Linear:} }E_{L}=\frac{1}{2}k_{l}\cdot x^{2}$ (2)
Torsional: $\text{{Torsional:} }E_{T}=\frac{1}{2}k_{\vartheta}\cdot\vartheta^{2}.$ (3)
@@
The individual stiffness and spring connectivity allow us to capture the limiting layer and dome inversion effects. We use a nonlinear spring to represent different dome mechanics (Fig. S1c), which are controlled by the parameters $k_{b}$, $\alpha$, and $d$. Linear and rotational springs are utilized to capture the limiting layer of the finger by combining stretching and bending energy (see Fig. 3a). The connection between the limiting layer and the nonlinear springs is modeled as an infinitely rigid connection, given that the strain is negligible compared to the rest of the lattice elements. By minimizing the sum of all contributions to the system ($E_{tot}=\sum_{i=1}^{n}E_{L}+E_{NL}+E_{T}$), the static problem is solved, and all possible finger configurations are predicted (see Fig. 3b).
@@
One of the primary challenges of spring-lattice models is accurately mapping the geometric parameters of the system to its stiffness constants. To address this, we integrate Finite Element (FE) simulations with Automatic Relevance Determination (ARD) regression [48] to derive expressions for $k_{b}$, $\alpha$, and $d$—the nonlinear spring parameters in Eq. 1—as functions of dome height ($H$), dome thickness ($t$), and dome curvature ($R$) (see Fig. S1). We considered different dimensionless relations ($\pi_{i}$) based on the most relevant geometric interactions affecting the dome’s energy [21]. These mechanics-informed interactions enhance the generality of the analysis by linking the mechanics to the unit’s geometry. For our model, we assume all spring constants can be expressed as functions of the shell’s load-carrying capacity ($\pi_{1}=\frac{t}{H}$), curvature-to-thickness ratio ($\pi_{2}=\frac{t}{R}$), and dome shallowness ($\frac{H}{R}$). Consequently, each parameter can be formulated as:
@@
Where $\mathbf{C}(\pi_{1},\pi_{2},\pi_{3})$ is a matrix of all the possible candidates and interactions of the non-dimensional relations, as shown in Eq. S.12. By using ARD, we can automatically determine the relevance of each feature in our model by examining the values of the hyperparameters, which determine the relevance of the weights $\xi_{\alpha}$, $\xi_{k}$, and $\xi_{d}$. As a result, we derive a general equation in terms of the three non-dimensional parameters (Eq. S.14, Eq. S.15, and Eq. S.16, see section A.2.3 for details). Using these derived expressions and a gradient-based optimization algorithm (see section A.2.4 for details), the final shape of the finger can be predicted by minimizing the system’s energy, which is now a function of the geometric parameters of the structure.
A comparison between FE simulations and our model for different stable states of the finger (see Fig. 3b) shows good agreement with the energy-based lattice model. The model is further extended to capture the structure’s dynamic response, allowing us to observe the material’s viscoelastic behavior and analyze metastable cases that energy minimization alone cannot predict. The dynamic response for different unit cell geometries is shown in Fig. S5b, and for different five-segmented fingers in Fig. S5c. The energy minimization approach and the dynamic model predict the same final state of the system after reaching a steady state. To explore the combined behavior of different dome geometries, we modeled a five-segment DPF topology with bistable and metastable units (see Fig. 3c). We stacked three bistable units with different dome heights ($H_{1}=4$ mm, $H_{2}=4.5$ mm, and $H_{3}=5$ mm) and two metastable units with the same dome height ($H_{4}=H_{5}=3$ mm). This configuration allows us to achieve three unique set points in the form of stable states and a gripper state that holds its shape for a given predictable time $\tau$ (see Fig. 3c). As shown in Fig. 3c, three different snap-through events and two snap-backs (dome-unit resets) are captured, adequately representing the mechanics of the combined phenomena.

### v2

@@ Abstract
Soft robots are distinguished by their flexibility and adaptability, allowing them to perform nearly impossible tasks for rigid robots. However, controlling their configuration is challenging due to their nonlinear material response and infinite deflection degrees of freedom. A potential solution to these challenges is to discretize the infinite-dimensional configuration space of soft robots into a finite but sufficiently large number of functional shapes. Leveraging this approach, we present a strategy for co-designing the desired tasks and morphology of pneumatically actuated soft robots with multiple encoded stable states. Our approach introduces a general method to capture the soft robots’ response using an energy-based analytical model, the parameters of which are obtained using Recursive Feature Elimination with lasso regression. The resulting lumped-parameter model facilitates inverse co-design of the robot’s morphology and planned tasks by embodying specific dynamics upon actuation and object interactions. The resulting computationally efficient model enables us to explore the configuration space in a tractable manner, allowing the inverse co-design of our robot morphology by setting desired set points with optimized stiffness. This strategy offers a framework for simplifying the control of soft robots by exploiting the nonlinear mechanics of multistable structures and embodying mechanical intelligence into soft material systems.
@@ 1 Introduction
Soft robots have gained interest due to their ability to interact with their environment, adapt to external stimuli, and protect against external disturbances [1, 2, 3]. The inherent safety provided by the characteristic low modulus of materials used in soft robots allows them to perform nearly impossible tasks for their rigid counterpart [4]. The interplay between soft mechanics and controls intrinsic to soft robotics gives rise to innovative solutions for tasks ranging from simple grasping to complex manipulation and object interaction [5, 6, 7]. However, this interplay also poses challenges complicating their modeling and control. These challenges include their infinite dimensionality, material nonlinearity, and large deformations that most soft robots exhibit [8]. As a result, sensory systems [9] and complex models/algorithms [10, 11] are required to implement open-loop control, leading to computationally demanding models to represent the robot’s behavior. Embedding sensing and control in the robots’ architecture [12] offers a new strategy to address some of the soft robotics’ main challenges, utilizing the robot morphology to reduce the use of complex algorithms and models and relying on its mechanical behavior. For example, granular material-based universal gripper [13] can adapt to and grasp different object shapes without needing closed-loop control based on their mechanical response. Zoe et al. [14] developed a methodology to use various soft actuators as pneumatic sensors, enabling distance sensing, shape sensing, profile scanning, and stiffness determination. Moreover, mechanical instabilities have been utilized for sequential programming and fast actuation. Yang et al. [15] explore the sequential programming of a soft gripper to perform grasping and twisting motion driven solely by mechanical instabilities in the gripper’s body. Yet another example by Lue Y. et al. [16] demonstrates the use of snap-through instabilities for fast actuation and reconfiguration. These examples illustrate how the intelligent co-design of the robot’s body morphology and functional behavior enables an approach to close the control loop for realizing desired tasks.
@@
Multistable structures offer an alternative path to achieve control of soft robotics with feedback via the programming of input-specific stable and defined shapes [17, 18, 19]. These types of structures display more than one energetically favorable configuration, giving the possibility to reach different final shapes from specific actuation inputs. They are often the result of assembling multiple classical bistable sub-structures, including constrained beams/trusses [20, 21], constrained dielectric elastomers [22], shells [23, 17, 24, 25], compliant mechanisms [26], or inflatable structures [27]. Combining multistable structures with soft robots can reduce the number of degrees of freedom and complexity to achieve control in soft systems, ultimately enhancing their functionality by providing the ability to program desired set points and sequences without closed-loop feedback [28, 29, 30], relying only on the mechanical behavior of the system [31]. Presently, a wide variety of soft machines leverage mechanical instabilities to improve the design and performance of their soft machines [32, 33, 34, 35] by programming different stable configurations and utilizing the rapid energy release during a snap-through instability [36, 37, 38, 39, 16]. Conrad et al. [40] show an electronics-free controller for a complex pneumatic system based on soft logic gates that can be incorporated into different soft machines. Similarly, Peretz et al. [41] utilized mechanical instabilities to create fluidic logic and reproduce valving functionality using entirely soft elements. Melancon et al. [42] used a bistable origami pattern to program different target points to its actuator, achieved by a single pressure source and predefined pressure path. Van Raemdonck et al. [43] geometrically tuned an actuator to generate complex actuation sequences from a single input by delaying the unit’s snap-through. Moreover, by leveraging the encoded stable states, the robot’s final configuration can be predicted based solely on its structural response. Consequently, the designed mechanical response defines the desired operational set points of the robot, all of which are stable. By transforming the infinite-dimensional deformation space into a finite number of stable states, the complexity of predicting the soft robot behavior is reduced, and its feedback can be reduced to a set of available states. This approach can simplify the complexity of predictive or data-driven algorithms currently needed to design and control soft robots, ultimately making soft robots more accessible and practical. However, the full potential of using multistability and programmed mechanical responses for soft robotics can only be realized by establishing computationally efficient methodologies for their morphological and functional co-design that circumvent the long runtime roadblocks imposed by the predominant use of conventional Finite Element (FE) packages to predict the geometry and response characteristics associated with the desired stable states.
@@
We introduce a framework for the modeling and inverse design of soft robotics with encoded mechanical behaviors by mapping into a discretized space the desired kinematic configurations leveraging mechanical instabilities (see Figure 1). We achieve this by incorporating different dome-shape programmable units addressable via pneumatic actuation featuring bistable, pseudo-bistable (metastable), and monostable [23, 44] mechanical behaviors (see Figure S2). As a result, different set points in the form of specific kinematical shapes and dynamic responses are embodied into the system morphology (see Figure 1), giving our soft robots the inherent capability of reaching different pre-programmed positions under open-loop inputs with almost zero error. This characteristic allows for robust soft robot design, as the bistable units provide intrinsic kinematic control dictated by the embodied mechanical response. The position control inherent to the designed multistable structure can serve as set points for triggering different control actions, decoupling the actuation from the robot’s final configuration (shape). The discretized response afforded by the proposed topology enables the introduction of an energy-based analytical modeling framework, allowing the efficient shape and structural property prediction of a multistable soft robot appendage topology (in our case, the Dome Phalanx Finger - DPF) that ultimately enables the inverse co-design of the robot’s functionality and control via the morphology. Our modeling approach is based on lumped parameter elements that can capture different mechanical behaviors of the dome unit via mechanics-informed identification, mapping the geometry of our robot to its mechanical response. The collective interaction of these elastic elements dictates the soft robot’s configuration, stiffness, and dynamic response. We demonstrate the versatility of our approach by performing the inverse co-design of different DPF topologies, giving specific target positions while maximizing the programmed states’ stiffnesses. Finally, we show that different dynamic behaviors can be encoded into the DPF morphology by leveraging the viscous effects of its constitutive material. We illustrate this new avenue for programming this class of robots’ time response through pre-programmed pick-and-place tasks. The presented modeling and inverse co-design approach leverages geometry and the resulting structural response to encode an electronics-free form of mechanical intelligence that simplifies the actuation and control of soft robots with implications on their accessibility, recyclability, security, and cost.
@@
Figure 1: Multistable soft robot with pre-programmed dynamics for positioning and stiffness. Our robot displays four grasping set points ($s_{0},s_{1},s_{3},s_{4}$) and embodied/morphological control. The robot’s morphology and mechanical behavior encodes the input and output dynamics in its energy landscape (G(s)). Feedback H(s) is derived from the dynamic response of the robot, whereby designed attractors for each stable set point are encoded in the robot geometry. The system’s output shows four set points as static minima in the energy landscape and attractors in the phase portrait, where $q$ and $\dot{q}$ are a schematic representation of the generalized coordinates of the robot. Set points (energy minima) are accessed by activating/inverting the dome units: State 1 $\rightarrow$ Stress-free state (State 1) and $\rightarrow$ Inverted state (State 2).
@@ 2.1 Embodied position and stiffness control
We begin by modifying the common PneuNet bending actuator topology [45] via the inclusion of dome-shaped units to create our Dome Phalanx Finger (DPF) topology (see Figure 2a). The finger derives its behavior from the mechanical response of the dome-shaped shell elements and the interaction between each segment. As the units invert, the top section of the finger expands, while the limiting layer (see Figure S1d) retains its original length. This results in global curvature and different final stable shapes, depending on the number of inverted dome units (see Figure S2b).
@@
These dome-shaped units program different mechanical responses and stable shapes into the system, effectively discretizing the infinite-dimensional deformation space into a manageable number of kinematic configurations dependent on the dome unit geometry, each attainable via well-defined, open-loop inputs. The mechanical response and stable configurations of the DPF can be tuned by adjusting the dome height per unit ($H_{i}$), dome thickness ($t$), number of segments, dome unit size (UC), pneumatic chamber thickness ($t_{\text{ch}}$), limiting layer thickness ($t_{\text{lim}}$), dome unit length ($U^{i}_{\text{L}}$), air chamber dimensions ($W_{\text{ch}}$ and $t_{\text{mid}}$), and the spacing between adjacent cells per unit ($U^{i}_{\text{sep}}$) (see Figure 2a and Figure S1a for reference). The geometric parameters of the dome units $H$ and $t$ dictate the mechanical response, determining whether it is monostable (Figure S1c(i)), metastable (Figure S1c(ii) and Movie 1), or bistable (Figure S1c(iii)). Furthermore, the dynamic response of the DPF can be controlled by the material properties and additional geometric parameters ($t_{\text{lim}}$, $t_{\text{ch}}$, $t_{\text{mid}}$), which significantly influence global curvature, system overshoot, and steady-state error (see Figure S3).
@@
Given the broad range of geometrical configurations, each finger can be geometrically tuned to exhibit different mechanical responses: a bistable response, where the final shape is retained after dome inversion [23, 17, 46] (Figure 2b $H_{i}$ = 4 mm and $H_{i}$ = 5 mm). A metastable response, where each unit undergoes snap-through after which its shape is retained for a time $t=\tau$, before returning to its stress-free state as a consequence of the combined effects of the unit’s viscoelastic material behavior and dome geometry [36, 47, 48] (Figure 2b, $H_{i}$ = 3 mm and Movie 1). Lastly, a monostable response, where the material behavior dominates over the geometric response and the finger responds similarly to a PneuNet bending actuator. As a result of the encoded mechanical responses, the actuation pressure required to invert the dome units is decoupled from the final shape, which is instead entirely determined by geometric parameters. Figure 2 d presents experimental measurements of the DPF tip displacement magnitude over time for various actuation pressures (dashed lines), demonstrating that the same final position is achieved despite different actuation magnitudes (Input pressure). Moreover, we leverage the material’s viscous response and structural properties to reset the system, whereby the designed unit metastability (dependent on dome height, material properties) and the duration of the applied load control the reset dynamics (see Figure 2c). The resulting freedom to design the DPF’s dynamics and set points allows for pre-programming a diverse set of tasks, sequences, and positions.
@@
Figure 2: Dome Phalanx Finger (DPF) with different encoded mechanical responses and behaviors. a) 5-segment DPF geometry parameters ($i=5$). b) Effect of dome height on finger response (Every segment with the same height). $H_{i}$ = 5 mm and $H_{i}$ = 4 mm show a bistable behavior, and $H_{i}$ = 3 mm shows a pseudo-bistable (metastable) response (self-resetting system). c) Bistable finger ($H_{i}=5$ mm) Tip displacement and pressure over time. d) Metastable finger ($H_{i}=3$ mm) Tip displacement and pressure over time.
e) Time lapse of two gripper architectures showing both bistable and pseudo-bistable responses.
@@
To showcase these unique capabilities, we developed a two DPF gripper architecture. By changing the dome height, we can program a gripper composed of two fingers to pick and hold an object after actuation (Figure 2c, $H_{i}$ = 4 mm) or pick an object and release it as the dome units reset (Figure 2c, $H_{i}$ = 3 mm), the latter of which is ideal for embodied pick-and-place task typically performed by conventional soft grippers. The number of stable states that can be encoded using the DPF morphology equals that of the number of constitutive units where each unit is tunable independently. Consequently, different unit geometries can be combined to achieve multiple stable states enabling a diversity of behaviors within the same overall geometry (Bistable + Metastable DPF). This allows for different energy minima and time responses to be embodied within the same finger, which can be accessed by a single pressure input and serve as discrete and functional robot configurations. To effectively design, combine, and predict the finger’s behavior based on its geometry, a simple yet reliable modeling framework is necessary. Given the broad design space and the computational cost of using FE models for our finger topology, we propose an energy-based spring model that captures the complex mechanical behaviors needed to design multistable soft robots with desired functional characteristics.
@@ 2.2 Model derivation and mechanics-informed parameter determination
Establishing an analytical model to capture the global behavior of the DPF requires consideration of both the unit cells’ and the global geometric parameters. By capturing the contribution of each constitutive unit of the DPF and their interactions, we can construct an energy landscape with a simplified model, where each of the minima corresponds to a stable state and a final position of the system. In this way, the state of the system ($x_{i}$) can be extracted from the resulting strain energy potential via a minimization process, whereby the DPF’s geometrical and stiffness characteristics dictates the programmed stable shapes. The DPF geometry is represented by a lattice array comprising nonlinear, linear, and torsional springs for the static response and point masses and dash-pots for the dynamic response (Figure 3a). The energy for the springs used in the lattice can be written as follows [49]:
Nonlinear: $\text{{Nonlinear:} }E_{NL}(x_{ij})=\frac{1}{2}k_{b}x_{ij}^{2}\left(1+\left((1-\alpha)\left(\left(\frac{x_{ij}}{d}\right)^{2}-2\frac{x_{ij}}{d}\right)\right)\right)$ (1)
Linear: $\text{{Linear:} }E_{L}(x_{ij})=\frac{1}{2}k_{l}\cdot x_{ij}^{2}$ (2)
Torsional: $\text{{Torsional:} }E_{T}(\vartheta)=\frac{1}{2}k_{\vartheta}\cdot\vartheta^{2}$ (3)
$E_{tot}(x_{ij})=\sum_{i=1}^{n}E_{L}(x_{ij})+E_{NL}(x_{ij})+E_{T}(x_{ij}).$ (4)
@@
where $x_{ij}=x_{i}-x_{j}$, $\vartheta$ is the angle between the rigid link and the limiting layer (see Figure S5a and S7c) and $k_{b}$, $k_{l}$, $k_{\vartheta}$ are stiffness parameters that depend on the DPF geometry. The individual stiffness and spring connectivity allow us to capture the limiting layer and dome inversion effects. We use a nonlinear spring to represent different dome mechanics (Figure S1c), which are described by the parameters $k_{b}$, $\alpha$, and $d$. Linear and rotational springs are utilized to capture the limiting layer of the finger by combining stretching and bending energies (see Figure 3a). The connection between the limiting layer and the nonlinear springs is modeled as an infinitely rigid connection, given that the strain is negligible compared to the rest of the lattice elements. By minimizing the sum of all contributions to the system (Equation 4), the static problem is solved, and all possible finger stable configurations are predicted (see Figure 3b).
@@
A primary challenge of spring-lattice models is to accurately map the geometric parameters of the system to its stiffness constants. To address this challenge, we integrate FE simulations (see section A.2) with Recursive Feature Elimination (RFE; see section A.3[50] and lasso regression to derive expressions for $k_{b}$, $\alpha$, and $d$—the nonlinear spring parameters in Equation 1—as functions of dome height ($H$), dome thickness ($t$), dome curvature ($R$) and Young Modulus ($E$) of a generic constitutive unit (see Figure S1). We considered different dimensionless relations ($\pi_{i}$) based on the most relevant geometric interactions affecting the dome’s energy [23]. These mechanics-informed interactions enhance the generality of the analysis by linking the mechanics to the unit’s geometry. For our model, we assume all spring constants can be expressed as functions of the shell’s load-carrying capacity ($\pi_{1}=\frac{t}{H}$), curvature-to-thickness ratio ($\pi_{2}=\frac{t}{R}$), and dome shallowness ($\pi_{3}=\frac{H}{R}$). Consequently, each parameter can be formulated as:
@@
where $\mathbf{C}(\pi_{1},\pi_{2},\pi_{3})$ is a matrix containing all possible candidates and interactions of the non-dimensional relations, as shown in Equation S.14. Using RFE, the relevance of each feature can be automatically determined by iteratively removing one feature at a time and observing the model’s coefficient of determination ($r^{2}$). This process yields the weights $\xi_{\alpha}$, $\xi_{k}$, and $\xi_{d}$ with fewer features while maintaining high accuracy. Notably, only the unit cell is used to determine the parameter features, and as a consequence, the DPF’s mechanical behavior is derived. A wide range of Young’s modulus ($E$), dome height-to-dome base ratio $\left(\frac{H}{r_{b}}\right)$, dome base radius ($r_{b}$), and thickness ($t$) are simulated to capture the entire design space of the DPF, including various soft materials ($5MPa\geq E\leq 40$ MPa). As a result, we derive a general equation in terms of the three non-dimensional parameters (Equation S.16, Equation S.17, and Equation S.18, see Section A.3.3 for details). Using these derived expressions and a gradient-based optimization algorithm (see Section A.4 for more information), the final shape of the finger can be predicted by minimizing the system’s energy, which is now expressed as a function of the geometric parameters of the structure.
Figure 3: An energy-based model for static and dynamic analysis of multistable soft robots based on the DPF topology. a) Spring lattice model with linear, rotational, and nonlinear springs. b) Static stable states predicted by the lattice model for $H_{1}$ = $H_{2}$ = 4 mm, $H_{3}$ = $H_{4}$ = 4.5 mm, and $H_{5}$ = 5 mm. c) Dynamic response for the soft robot with four bistable units and two metastable units ($H_{1}=H_{2}=H_{3}=H_{4}=5$ mm, and $H_{5}=H_{6}=3$ mm). The model captures the snap-through of each bistable unit and the snap-back of the metastable units (See Movie 2).
A comparison between FE simulations and the proposed model for different stable states of the finger (see Figure 3b) and different design cases (see Figure S10) shows good agreement with the energy-based lattice model, while reducing the computational time by more than $10^{-3}\times$ (see Table S2). The model is further extended to capture the structure’s dynamic response, allowing us to observe the material’s viscoelastic behavior and analyze metastable cases that energy minimization alone cannot predict. The dynamic response for different unit cell geometries is shown in Figure S7b, and for different five-segmented fingers in Figure S7d, where the dynamic model predicts the same final state of the system after reaching a steady state. To explore the combined behavior of different dome geometries, we modeled a five-segment DPF topology with bistable and metastable units (see Figure 3c). We stacked four bistable units ($H_{1}=H_{2}=H_{3}=H_{4}=5$ mm) and two metastable units with the same dome height ($H_{5}=H_{6}=3$ mm). This configuration allows us to achieve one unique set point: a stable state and a time-dependent state that holds its shape for a given predictable time $\tau$ (see Figure 3c). The resulting dynamic response features five different snap-through events and two snap-backs (dome-unit resets), adequately representing the complex dynamics displayed by the DPF (Figure 3c).